# ReGNet: Reciprocal Space-Aware Long-Range Modeling for Crystalline Property Prediction

### v2

@@ 3.1 Representation and Embedding
Invariant crystal graph and embedding.
@@
=$\mathcal{E}=\{e_{ij}:\|\bm{p}_{i}-\bm{p}_{j}\|_{2}\leq r_{\text{cut}},\ \forall i,j\in V\},$
@@
$h^{0}_{\text{global}}=\sigma(W_{r}h^{0}_{\text{local}}),$
where $W_{r}$ is a learnable weight matrix. We use ${h}^{0}_{\text{global}}$ as initial long-range node features. The invariant edge feature $||\mathbf{p}_{i}-\mathbf{p}_{j}||_{2}$ is defined by the Euclidean distance between node pair $i$ and $j$, are scaled by $c/||\mathbf{p}_{i}-\mathbf{p}_{j}||_{2}$, where c is a chosen constant to mimic the pairwise potential in Lin et al. (2023). These values are then embedded using radial basis function (RBF) kernels to get the initial edge features ${v}_{ij}^{e}$. More details are shown in Section B.2
@@
To effectively extract features from long-range information in reciprocal space, we propose using fractional coordinates $\bm{f}$ and basis reciprocal lattice vectors $\bm{k}_{m}$ with continuous filters convolution. Together, they enable a precise and symmetry-consistent representation of periodic structures (discussed in Section 2). Our approach differs from previous methods (Kosmala et al., 2023) for the following reasons. First, for crystalline materials, Cartesian coordinates can introduce ambiguities, as identical crystal structures may be expressed differently due to periodic transformations, such as translations or rotations. For example, the lattice matrices $\textbf{L}=[\bm{\ell}_{1},\bm{\ell}_{2},\bm{\ell}_{3}]^{\top}$ and $\textbf{L}^{\prime}=[\bm{\ell}_{1}+\bm{\ell}_{2},\bm{\ell}_{2},\bm{\ell}_{3}]^{\top}$ describe identical periodic patterns but differ in their Cartesian representations. In contrast, fractional coordinates normalize atomic positions relative to the unit cell, ensuring invariance under periodic transformations and eliminating redundancies in Cartesian representations. Second, it relies on a predefined grid size as hyperparameters to cover all relevant frequencies as $\bm{k}$, which could disrupt the unit cell symmetry and substantially increase computational costs, particularly for large unit cells and complex systems. Instead, our method leverages the natural periodicity of the lattice in materials, avoiding redundant grid parameterization and ensuring a symmetry-consistent representation of long-range interactions. Therefore, by combining fractional coordinates $\bm{f}$ and basis reciprocal lattice vectors $\bm{k}_{m}$, we achieve an effective framework for feature extraction in reciprocal space, accurately capturing the long-range information of crystal materials.
With the aforementioned representations, our model enables both short-range and long-range message passing to effectively capture comprehensive material structural features.
@@ 3.2 Short-Range Message Passing
We employ the short-range block in Figure 2(b) to update the atomic representations using an invariant graph neural network with multiple message-passing layers. Each node aggregates information only from its geometric neighbors within the cutoff radius $r_{\text{cut}}$. Therefore, we naturally interpret this as capturing the short-range information in material representations. The computational process at the $\ell$-th message-passing layer for a node $i$ is expressed as:
@@ 3.3 Long-Range Message Passing
To complement the short-range modeling, we employ the long-range block in Figure 2(b) to globally update the atomic representations using the ReciprocalBlock. Specifically, this block leverages Fourier transform-based continuous filters with neural networks to encode periodic structures from reciprocal space. Besides, as discussed in Section 3.1, we utilize atomic fractional coordinates $\bm{f}$ and basis reciprocal lattice vectors $\bm{k}_{m}$ to extract structural information from reciprocal space and update the node features in the model block. The global embeddings are computed iteratively:
$h^{\ell+1}_{\text{global}}=\text{ReciprocalBlock}({h}_{\text{global}}^{\ell},\bm{f},\bm{k}_{m}),$
@@
${r}_{m}=\sum_{j\in\mathcal{I}_{m}}{h}_{j,\text{global}}^{\ell}\cdot\exp(-i\bm{k}_{m}^{\top}\bm{f}_{j}),$
@@
Subsequently, the reciprocal embedding ${r}_{m}$ undergoes an inverse Fourier transform to be mapped back into real space for the atoms in material $m$. This process can be expressed as below:
$\tilde{h}_{\text{global}}^{\ell}=\sum_{j\in\mathcal{I}_{m}}\exp(i\bm{k}_{m}^{\top}\bm{f}_{j})\cdot{r}_{m}\cdot\mathbf{W}_{\text{filter}},$
where $\exp(i\bm{k}_{m}^{\top}\bm{f}_{j})$ performs the inverse Fourier transform to map the reciprocal representation back to the real domain. $\mathbf{W}_{\text{filter}}$ is a trainable filter refining the reciprocal embedding, selectively emphasizing important contributions from the reciprocal space features to model long-range interactions. In addition, the prefactor $\frac{1}{\Omega}$ in eq. 2 is incorporated into the learned filter $\mathbf{W}_{\text{filter}}$, without requiring explicit normalization by the system volume. In a word, this operation effectively incorporates long-range interactions and periodicity from the reciprocal domain into the real-space representation of each atom $j\in\mathcal{I}_{m}$.
@@
${h}^{\ell+1}_{\text{global}}={h}^{\ell}_{\text{global}}+\tilde{{h}}^{\ell}_{\text{global}}.$
@@
3.4 Hierarchical Embedding Integration
Our model employs multiple ReGNet blocks as message passing layers to integrate local and global information at each layer. Within each block, node embeddings are updated by combining contributions from short-range graph message passing and long-range reciprocal-based message passing. The resulting integrated embeddings, formed by adding up both embedding updates during each message-passing step, are propagated to the next block. This enables the model to iteratively refine hierarchical representations of material structures. This design facilitates the effective capture of both local atomic-scale features and global lattice-scale features throughout the network.
3.5 Decoder Block
After completing the iterative message-passing steps, node features are aggregated within each graph using mean pooling, which captures the overall structural information of the material. The resulting graph-level representation is then processed through fully connected layers to predict the target material property. While our novel reciprocal space representations effectively captures long-range interactions and global periodicity, it still predicts properties individually Schütt et al. (2017); Choudhary & DeCost (2021); Yan et al. (2024). Therefore, for higher efficiency, we employ an MTL decoder for multi-property prediction, consisting of MoE layers and task-specific heads. More details can be found in Appendix C.
4 Related Work
Crystal property prediction.
Crystal property prediction has been widely studied using both physics-based and deep-learning approaches. Traditional physics-based methods, such as Coulomb matrices (Rupp et al., 2012; Elton et al., 2018), effectively model ionic and metallic materials but lack generalization due to constraints like permutation invariance. Deep learning has introduced more flexible predictions by representing crystals as chemical formulas and using sequence models (Jha et al., 2018; Goodall & Lee, 2020; Wang et al., 2021). More recent methods leverage 3D geometric structures, modeling crystals as 3D graphs with graph neural networks (GNNs) (Schütt et al., 2017; Gasteiger et al., 2020). Key advances include CGCNN (Xie & Grossman, 2018), which utilized multi-edge graphs to model periodic invariance, inspiring methods such as ALIGNN (Choudhary & DeCost, 2021), which incorporated angle-based features, and Matformer (Yan et al., 2022), which captured periodic patterns using self-connecting edges. PotNet (Lin et al., 2023) proposed physics-informed edge features that embed the infinite summation of pre-defined interatomic potentials, while Crystalformer (Taniai et al., 2024) encoded periodic structures with infinitely connected attention. Yan et al. (2024) introduced iComFormer with invariant descriptors and eComFormer with equivariant vectors. Despite significant advancements, many approaches still face limitations in accurately capturing periodic patterns and long-range interactions.
Long-range interaction modeling.
Modeling long-range interactions remains a significant challenge. Early approaches augmented physical equations with hand-crafted terms (Staacke et al., 2021) or predicted atomic charges (Unke et al., 2021), while recent methods focus on data-driven learning. Yu et al. (2022) computed scalar-valued structure factors and integrates them with GNN outputs only in the final layer, limiting interaction between reciprocal space and local features. In contrast, Kosmala et al. (2023) integrated structure factor embeddings directly into GNN message passing, enabling feedback to node embeddings. However, it requires extensive hyperparameter tuning for grid sizes ($N_{x},N_{y},N_{z}$) to define the supercell, which can disrupt the symmetry of the unit cell, alter its space group, and substantially increase computational costs, particularly for systems with complex symmetries. Lin et al. (2023) considered the infinite summations of pairwise atomic interactions with additional computational cost $\mathcal{O}(n^{2})$, relied on precomputed, non-trainable physical potentials as edge features, which limited adaptability to diverse material systems and further increased computational overhead.
Reciprocal space.
@@ 4 Related Work
CGCNN 31 0.292 0.047 0.077
SchNet 33 0.345 0.066 0.099
MEGNET 30 0.307 0.060 0.099
GATGNN 33 0.280 0.045 0.075
ALIGNN 22 0.218 0.051 0.078
Matformer 21.0 0.211 0.043 0.073
PotNet 18.8 0.204 0.040 0.0650
Crystalformer 18.6 0.198 0.0377 0.0689
eComFormer 18.16 0.202 0.0417 0.0729
iComFormer 18.26 0.193 0.0380 0.0637
ReGNet 17.07 0.189 0.0328 0.0628
While (Kosmala et al., 2023) incorporates reciprocal space modeling, their approach is tailored to molecular systems and evaluated on semi-periodic surface structures from the OC20 dataset, which includes relaxed small-molecule adsorbates (e.g., CO, H2O, NH3) on metal surfaces (e.g., Pt, Cu, Ni). These systems lack full 3D periodicity and space group symmetries typical of bulk crystals. As a result, the study does not assess reciprocal space modeling in fully periodic crystalline materials such as perovskites (e.g., SrTiO3) or semiconductors (e.g., GaAs), which are typical and important materials. In contrast, our method is explicitly designed for bulk crystals and directly encodes their symmetry and periodicity. Separately, Taniai et al. (2024) introduces reciprocal space via fixed analytical Fourier-Gaussian formulations within a Transformer-style architecture. However, their reciprocal-space component is non-trainable and presented only as an optional variant in the appendix. Our proposed ReGNet instead employs fully learnable reciprocal-space filters, trained end-to-end to capture long-range interactions beyond the capacity of fixed-function decays.
5 Experimental Studies
We evaluate ReGNet on three widely recognized datasets in materials science, Materials Project (Chen et al., 2019), JARVIS-DFT (Choudhary et al., 2020), and MatBench (Dunn et al., 2020). Through extensive experiments, ReGNet demonstrates strong performance across a wide range of crystal properties, highlighting its effectiveness in modeling crystalline materials with reciprocal space.
For JARVIS and MP, to ensure consistency, we adopt experimental settings aligned with prior works, including Matformer (Yan et al., 2022) and PotNet (Lin et al., 2023). These datasets cover a range of scales, including 69239 crystals for large-scale tasks, 18171 crystals for medium-scale tasks, and 5450 crystals for small-scale tasks. To further assess performance, we evaluate ReGNet on the MatBench tasks e_form (132752 crystals) and jdft2d (636 crystals), following the experimental setup in (Yan et al., 2024). We benchmark against several baseline methods, including CGCNN (Xie & Grossman, 2018), SchNet (Schütt et al., 2017), MEGNet (Chen et al., 2019), GATGNN (Louis et al., 2020), ALIGNN (Choudhary & DeCost, 2021), Matformer (Yan et al., 2022), PotNet (Lin et al., 2023), Crystalformer (Taniai et al., 2024), ComFormer (Yan et al., 2024), MODNet (De Breuck et al., 2021) , coGN (Ruff et al., 2024), and M3GNet (Chen & Ong, 2022). All experiments are conducted on NVIDIA A100 GPUs. Additional details on the experimental configurations and implementation settings are provided in Appendix B.
5.1 Experimental Results
The Materials Project (MP)
On the Materials Project dataset, ReGNet consistently outperforms prior methods across all four benchmark tasks, achieving state-of-the-art results with notable improvements. As shown in table 1, ReGNet achieves lower MAE than iComFormer on all target properties, including a 6.5% reduction in formation energy error, 2.1% in bandgap, 13.7% in bulk modulus, and 1.4% in shear modulus.
JARVIS-DFT
As shown in Table 2, ReGNet surpasses baseline models in formation energy, $E_{\text{total}}$, and band gap (MBJ), while achieving competitive results on $E_{\text{hull}}$ and band gap (OPT). Notably, ReGNet consistently surpasses PotNet across all tasks and improves upon iComFormer with a 7.7% lower MAE on band gap (MBJ) and a 6.9% reduction on total energy prediction.
MatBench
Experimental results on the MatBench benchmark are summarized in Table 3, reported in terms of MAE and RMSE with standard deviations over five runs. ReGNet achieves state-of-the-art performance on both the large-scale e_form task (over 130K crystals) and the small-scale jdft2d task (636 crystals), demonstrating strong generalization across both large and limited training data. Notably, it outperforms models that incorporate geometric features such as bond angles, including ALIGNN and M3GNet, as well as coGN, which employs a connectivity-optimized nested graph representation. ReGNet also surpasses MODNet, a leading model based on physical descriptors with feature selection. These results highlight the effectiveness of incorporating reciprocal space information into crystal representations.
@@ 5.1 Experimental Results
CGCNN 63 0.20 78 0.41 170
SchNet 45 0.19 47 0.43 140
MEGNET 47 0.145 58 0.34 84
GATGNN 47 0.170 56 0.51 120
ALIGNN 33.1 0.142 37 0.31 76
Matformer 32.5 0.137 35 0.30 64
PotNet 29.4 0.127 32 0.27 55
Crystalformer 30.6 0.128 32 0.27 46
eComFormer 28.4 0.124 32 0.28 44
iComFormer 27.2 0.122 28.8 0.26 47
ReGNet 27.0 0.126 26.8 0.24 43
@@
MODNet 44.8 $\pm$ 3.9 88.8 $\pm$ 7.5 33.2 $\pm$ 7.3 96.7 $\pm$ 40.4
ALIGNN 21.5 $\pm$ 0.5 55.4 $\pm$ 5.5 43.4 $\pm$ 8.9 117.4 $\pm$ 42.9
coGN 17.0 $\pm$ 0.3 48.3 $\pm$ 5.9 37.2 $\pm$ 13.7 101.2 $\pm$ 55.0
M3GNet 19.5 $\pm$ 0.2 - 50.1 $\pm$ 11.9 -
eComFormer 16.5 $\pm$ 0.3 45.4 $\pm$ 4.7 37.8 $\pm$ 9.0 102.2 $\pm$ 46.4
iComFormer 16.5 $\pm$ 0.3 43.2 $\pm$ 3.7 34.8 $\pm$ 9.9 96.1 $\pm$ 46.3
ReGNet 16.1 $\pm$ 0.2 42.9 $\pm$ 3.5 31.6 $\pm$ 6.3 75.6 $\pm$ 34.7
Overall, ReGNet delivers state-of-the-art performance across various crystal property prediction tasks on three widely used crystal benchmarks. The superior performance of ReGNet can be attributed to several key factors. Notably, ReGNet leverages trainable Fourier filters from reciprocal space to capture long-range interactions while preserving the periodic symmetry of crystal structures, which are fundamental to the accurate and physically grounded representation of crystalline materials. In contrast, prior methods rely on fixed physical forms or predefined grid sizes that require tuning. Besides, by integrating reciprocal-space-based global information with GNN-driven local modeling, ReGNet provides a comprehensive representation of crystalline materials.

### v3

@@ 4.1 Representation and Embedding
Crystal graph and embedding.
@@
$\mathcal{E}=\{e_{ij}:\|\bm{p}_{i}-\bm{p}_{j}\|_{2}\leq r_{\text{cut}},\ \forall i,j\in V\},$ (4)
@@
$h^{0}_{\text{global}}=\sigma(W_{r}h^{0}_{\text{local}}),$ (5)
[…]
@@
To effectively extract features from long-range information in reciprocal space, we propose using fractional coordinates $\bm{f}$ and reciprocal lattice vectors $\bm{k}$ (Section˜3) with continuous filters convolution. This approach differs from previous methods for the following reasons. First, for crystalline materials, Cartesian coordinates can introduce ambiguities, as identical crystal structures may be expressed differently due to periodic transformations, such as translations or rotations. For example, the lattice matrices $\textbf{L}=[\bm{\ell}_{1},\bm{\ell}_{2},\bm{\ell}_{3}]^{\top}$ and $\textbf{L}^{\prime}=[\bm{\ell}_{1}+\bm{\ell}_{2},\bm{\ell}_{2},\bm{\ell}_{3}]^{\top}$ describe identical periodic patterns but differ in their Cartesian representations. In contrast, fractional coordinates normalize atomic positions relative to the unit cell, ensuring consistent representation under periodic transformations and eliminating redundancies in Cartesian representations. Second, Kosmala et al. (2023) relies on a predefined grid size as hyperparameter to cover all relevant frequencies, which could disrupt the unit cell symmetry and substantially increase computational costs, particularly for large unit cells and complex systems. Instead, by combining fractional coordinates $\bm{f}$ and basis reciprocal lattice vectors $\bm{k}_{m}$, we achieve an effective framework for feature extraction in reciprocal space, accurately capturing the long-range information of crystals.
With the aforementioned representations of crystalline materials, our model enables both short-range and long-range message passing to effectively capture comprehensive structural features.
@@ 4.2 Short-Range Message Passing
We employ the short-range block in Figure˜2 to update the atomic representations using the graph neural network with multiple message-passing layers. Each node aggregates information only from its geometric neighbors within the cutoff radius $r_{\text{cut}}$. Therefore, we naturally interpret this as capturing the short-range information in material representations. The computational process at the $\ell$-th message-passing layer for a node $i$ is expressed as:
@@ 4.3 Long-Range Message Passing
To complement the short-range modeling, we employ the long-range block in Figure˜2 to globally update the atomic representations using the ReciprocalBlock. Specifically, this block leverages Fourier transform-based continuous filters with neural networks to encode periodic structures from reciprocal space. Besides, as discussed in Section˜4.1, we utilize atomic fractional coordinates $\bm{f}$ and basis reciprocal lattice vectors $\bm{k}_{m}$ within material $m$ to extract structural information from reciprocal space and update the node features in the model block. The global embeddings are computed iteratively:
$h^{\ell+1}_{\text{global}}=\text{ReciprocalBlock}({h}_{\text{global}}^{\ell},\bm{f},\bm{k}_{m}),$ (7)
@@
${r}_{m}=\sum_{j\in\mathcal{I}_{m}}{h}_{j,\text{global}}^{\ell}\cdot\exp(-i\bm{k}_{m}^{\top}\bm{f}_{j}),$ (8)
@@
Subsequently, the reciprocal embedding ${r}_{m}$ undergoes an inverse Fourier transform to be mapped back into real space for the atoms in material $m$. This process can be expressed as follows:
$\tilde{h}_{\text{global}}^{\ell}=\sum_{j\in\mathcal{I}_{m}}\exp(i\bm{k}_{m}^{\top}\bm{f}_{j})\cdot{r}_{m}\cdot\mathbf{W}_{\text{filter}},$ (9)
where $\exp(i\bm{k}_{m}^{\top}\bm{f}_{j})$ performs the inverse Fourier transform to map the reciprocal representation back to the real domain. $\mathbf{W}_{\text{filter}}$ is a trainable filter refining the reciprocal embedding, selectively emphasizing important contributions from the reciprocal space features to model long-range interactions. The prefactor $\frac{1}{\Omega}$ in eq.˜3 is incorporated into the learned filter $\mathbf{W}_{\text{filter}}$, without requiring explicit normalization by the system volume. In a word, this operation effectively incorporates long-range interactions from the reciprocal domain into the real space representation of each atom $j\in\mathcal{I}_{m}$.
@@
${h}^{\ell+1}_{\text{global}}={h}^{\ell}_{\text{global}}+\tilde{{h}}^{\ell}_{\text{global}}.$ (10)
@@
CGCNN (Xie & Grossman, 2018) 31 0.292 0.047 0.077
SchNet (Schütt et al., 2017) 33 0.345 0.066 0.099
MEGNET (Chen et al., 2019) 30 0.307 0.060 0.099
GATGNN (Louis et al., 2020) 33 0.280 0.045 0.075
ALIGNN (Choudhary & DeCost, 2021) 22 0.218 0.051 0.078
Matformer (Yan et al., 2022) 21.0 0.211 0.043 0.073
PotNet (Lin et al., 2023) 18.8 0.204 0.040 0.0650
Crystalformer (Taniai et al., 2024) 18.6 0.198 0.0377 0.0689
eComFormer (Yan et al., 2024) 18.16 0.202 0.0417 0.0729
iComFormer (Yan et al., 2024) 18.26 0.193 0.0380 0.0637
CrystalFramer (Ito et al., 2025) 17.2 0.185 0.0338 0.0677
ReciNet 17.07 0.189 0.0328 0.0628
4.4 Hierarchical Embedding Integration
Our model employs multiple ReciNet blocks as message passing layers to integrate local and global information at each layer. Within each block, node embeddings are updated by combining contributions from short-range graph message passing and long-range reciprocal-based message passing. The resulting integrated embeddings, formed by adding up both embedding updates during each message-passing step, are propagated to the next block. This enables the model to iteratively refine hierarchical representations of material structures. This design facilitates the effective capture of both local atomic-scale features and global lattice-scale features throughout the network.
4.5 Decoder Block
After completing the iterative message-passing steps, node features are aggregated within each graph using mean pooling, which captures the overall structural information of the material. The resulting graph-level representation is then processed through fully connected layers to predict the material property. While our novel reciprocal space representation effectively captures long-range interactions and global periodicity, it still predicts properties individually (Schütt et al., 2017; Choudhary & DeCost, 2021; Yan et al., 2024). To validate the representation ability of the ReciNet encoder, we also explore its application in a multi-task learning (MTL) setting. This variant, ReciNet-MT, couples the shared encoder with an efficient mixture-of-experts (MoE) decoder to predict multiple properties simultaneously, as shown in Figure˜2. More details can be found in Appendix˜C.
@@ 4.5 Decoder Block
CGCNN (Xie & Grossman, 2018) 63 0.20 78 0.41 170
SchNet (Schütt et al., 2017) 45 0.19 47 0.43 140
MEGNET (Chen et al., 2019) 47 0.145 58 0.34 84
GATGNN (Louis et al., 2020) 47 0.170 56 0.51 120
ALIGNN (Choudhary & DeCost, 2021) 33.1 0.142 37 0.31 76
Matformer (Yan et al., 2022) 32.5 0.137 35 0.30 64
PotNet (Lin et al., 2023) 29.4 0.127 32 0.27 55
Crystalformer (Taniai et al., 2024) 30.6 0.128 32 0.27 46
eComFormer (Yan et al., 2024) 28.4 0.124 32 0.28 44
iComFormer (Yan et al., 2024) 27.2 0.122 28.8 0.26 47
ReciNet 27.0 0.126 26.8 0.24 43
5 Experimental Studies
We evaluate ReciNet on three widely recognized datasets in materials science, Materials Project (Chen et al., 2019), JARVIS-DFT (Choudhary et al., 2020), and MatBench (Dunn et al., 2020). Through extensive experiments, ReciNet demonstrates strong performance across various crystal properties, highlighting its effectiveness in modeling crystalline materials with reciprocal space.
For JARVIS and MP, to ensure consistency, we adopt experimental settings aligned with prior works, including Matformer (Yan et al., 2022) and PotNet (Lin et al., 2023). These datasets cover a range of scales, including 69,239 crystals for large-scale tasks, 18,171 crystals for medium-scale tasks, and 5,450 crystals for small-scale tasks. To further assess performance, we evaluate ReciNet on the MatBench tasks e_form (132,752 crystals) and jdft2d (636 crystals), following the experimental setup in Yan et al. (2024). We benchmark against major baseline methods, including CGCNN (Xie & Grossman, 2018), SchNet (Schütt et al., 2017), MEGNet (Chen et al., 2019), GATGNN (Louis et al., 2020), ALIGNN (Choudhary & DeCost, 2021), Matformer (Yan et al., 2022), PotNet (Lin et al., 2023), Crystalformer (Taniai et al., 2024), ComFormer (Yan et al., 2024), CrystalFramer (Ito et al., 2025), MODNet (De Breuck et al., 2021) , coGN (Ruff et al., 2024), and M3GNet (Chen & Ong, 2022). We exclude CrystalFormer from the JARVIS comparison to ensure fairness, as it requires significantly longer training time than other baselines, as shown in Table˜5. We also compare ReciNet with reciprocal space related methods, PotNet (Lin et al., 2023), EwaldMP (Kosmala et al., 2023), and Crystalformer (Taniai et al., 2024) variant, to highlight the advantage of ReciNet. Experiments are conducted on NVIDIA A100 GPUs. Additional details on the experimental configurations and implementation settings are provided in Appendix B.
5.1 Experimental Results
The Materials Project (MP). On the Materials Project dataset, ReciNet consistently outperforms prior methods across all four benchmark tasks, achieving state-of-the-art results with notable improvements. As shown in table˜1, ReciNet achieves lower MAE than iComFormer on all target properties, including a 6.5% reduction in formation energy error, 2.1% in bandgap, and 13.7% in bulk modulus.
JARVIS-DFT.
As shown in Table˜2, ReciNet surpasses baseline models in formation energy, $E_{\text{total}}$, and band gap (MBJ), while achieving competitive results on $E_{\text{hull}}$ and band gap (OPT). Notably, ReciNet consistently surpasses PotNet across all tasks and improves upon iComFormer with a 7.7% lower MAE on band gap (MBJ) and a 6.9% reduction on total energy prediction.
MatBench.
Experimental results on the MatBench benchmark are summarized in Table˜3, reporting MAE and RMSE with standard deviations over five runs. ReciNet achieves SOTA performance on both the large-scale e_form task (over 130K crystals) and the small-scale jdft2d task (636 crystals), demonstrating strong generalization across both large and limited training data. It outperforms models that incorporate geometric features such as bond angles, including ALIGNN and M3GNet, as well as coGN, which employs a connectivity-optimized nested graph representation. ReciNet also surpasses MODNet, a model based on physical descriptors with feature selection. These results highlight the effectiveness of incorporating reciprocal space information into crystal representations.
@@ 5.1 Experimental Results
MODNet (De Breuck et al., 2021) 44.8 $\pm$ 3.9 88.8 $\pm$ 7.5 33.2 $\pm$ 7.3 96.7 $\pm$ 40.4
ALIGNN (Choudhary & DeCost, 2021) 21.5 $\pm$ 0.5 55.4 $\pm$ 5.5 43.4 $\pm$ 8.9 117.4 $\pm$ 42.9
coGN (Ruff et al., 2024) 17.0 $\pm$ 0.3 48.3 $\pm$ 5.9 37.2 $\pm$ 13.7 101.2 $\pm$ 55.0
M3GNet (Chen & Ong, 2022) 19.5 $\pm$ 0.2 - 50.1 $\pm$ 11.9 -
eComFormer (Yan et al., 2024) 16.5 $\pm$ 0.3 45.4 $\pm$ 4.7 37.8 $\pm$ 9.0 102.2 $\pm$ 46.4
iComFormer (Yan et al., 2024) 16.5 $\pm$ 0.3 43.2 $\pm$ 3.7 34.8 $\pm$ 9.9 96.1 $\pm$ 46.3
ReciNet 16.1 $\pm$ 0.2 42.9 $\pm$ 3.5 31.6 $\pm$ 6.3 75.6 $\pm$ 34.7
Reciprocal space comparison.
To further demonstrate the superiority of reciprocal space design in ReciNet, we provide detailed comparisons with machine learning models that incorporate reciprocal space information, including PotNet, EwaldMP, and Crystalformer, as shown in Table˜12. Detailed results and discussions are presented in Appendix B.6. In summary, ReciNet consistently achieves lower MAE across all five crystal properties on the JARVIS dataset compared to these baselines.
Overall, ReciNet delivers state-of-the-art performance across various crystal properties on three widely used crystal benchmarks. This success is rooted in an architectural design that captures the full hierarchy of physical interactions in crystals. ReciNet overcomes the locality bias by operating in reciprocal space, the natural domain for periodic systems. Specifically, the usage of fractional coordinates $f$ and reciprocal lattice vectors $k$ constructs a Fourier series representation of the crystal. This process adaptively learns long-range representation while reserving crystal-naive symmetry.
Table 4: Ablation studies. Reciprocal space and model depth in test MAE on the JARVIS dataset.
Method Reciprocal Blocks Form. Energy Bandgap(OPT) $E_{total}$ Bandgap(MBJ) $E_{hull}$ ReciNet w/o Reciprocal 3 29.6 0.135 32.5 0.295 68.8 ReciNet w/o Reciprocal 4 28.9 0.132 31.0 0.290 65.6 ReciNet 3 27.8 0.127 27.2 0.242 52.9 ReciNet 4 27.0 0.126 26.8 0.239 48.0